PERIODIC AUTOMORPHISMS OF SURFACES: INVARIANT CIRCLES AND MAXIMAL ORDERS

HANSJÖRG GEIGES AND DIEGO RATTAGGI

Abstract. W. H. Meeks has asked the following question: For what g does every (orientation preserving) periodic automorphism of a closed orientable surface of genus g have an invariant circle? A variant of this question due to R. D. Edwards asks for the existence of invariant essential circles. Using a construction of Meeks we show that the answer to his question is negative for all but 43 values of g ≤ 10000, all of which lie below g = 105. We then show that the work of S. C. Wang on Edwards’ question generalizes to nonorientable surfaces and automorphisms of odd order. Motivated by this, we ask for the maximal odd order of a periodic automorphism of a given nonorientable surface. We obtain a fairly complete answer to this question and also observe an amusing relation between this order and Fermat primes.